OpenStudy (anonymous):

How do I find the surface area and volume of a cube with an equation of 4x-5?

OpenStudy (anonymous):

v=s^3 v=(4x-5)^3 A=6(4X-5)^2

OpenStudy (anonymous):

Thanks